OpenStudy (anonymous):

please help! A+B+C=180°, cos B + cos C ?? a. 2 cos (A/2) cos {(B-C)/2} b. 2 sin (A/2) cos {(B-C)/2} c. 2 cos (C/2) cos {(B-C)/2} d. 2 cos (B/2) cos (C/2) e. 2 sin (B/2) sin (C/2) (choose one answer that is correct and please give the explanation)

OpenStudy (mayankdevnani):

its option B

OpenStudy (mayankdevnani):

Here its explanation :- \[\huge \bf Given: A+B+C=180^o\] so \[\huge \bf B+C=180-A\] \[\huge \bf \frac{B+C}{2}=\frac{180}{2}-\frac{A}{2}\] Remember this formula :- \[\huge \bf Cos~B+Cos~C=2\cos \frac{B+C}{2}\cos \frac{B-C}{2}\] here, \[\huge \bf \frac{B+C}{2}=90^o-\frac{A}{2}\] therefore,we get \[\huge \bf 2\cos (90-\frac{A}{2})\cos \frac{B-C}{2}\] we know that \(\large \bf \cos(90- \theta)=\sin \theta\) here in this case ,we get \[\huge \bf 2\sin \frac{A}{2}\cos \frac{B-C}{2}\]

OpenStudy (mayankdevnani):

therefore,option B is correct

OpenStudy (mayankdevnani):

*remember this formula:- \[\large \bf \cos~B+\cos~C=2\cos \frac{B+C}{2}\cos \frac{B-C}{2}\]

OpenStudy (mayankdevnani):

@gabriellalyn understood?

OpenStudy (anonymous):

yea i got it. thank you!! @mayankdevnani

OpenStudy (mayankdevnani):

welcome :)